When a 25000-kg fighter airplane lands on the deck of the aircraft carrier, the carrier sinks 0.23cm deeper into the water.

a) Determine the cross-sectional area of the carrier. Assume the density of seawater is 1025 kg/m3.

A = ???
Answer #1


Density = mass/volume

volume = cross sectional area * deep of the water

Density = mass/(cross sectional area * deep of the water)

cross sectional area = mass/(density*deep of the water)

cross sectional area = 25000/(1025*0.23*10^-2) = 10604.45 m^2